# Fixed-point Inversion for Text-to-image diffusion models

### v1

@@ 6 Additional Qualitative Results
Figure 14: Generating rare concepts based on a few examples with the method of [18] that heavily depends on the diffusion-inversion quality. In our comparison, we evaluate FPI alongside DDIM (refer to the discussion in the main paper). DDIM frequently struggles to produce a realistic representation of certain objects (such as Oxygen-mask or Patas Monkey) or an accurate depiction of specific concepts (like Tiger-cat or Pay-phone). However, the use of FPI rectifies these issues in the results. For a detailed quantitative comparison, please refer to the main paper. See the main paper for a quantitative comparison.
Figure Fixed-point Inversion for Text-to-image diffusion models compares real image editing. While other methods struggle to achieve proper editing or take a considerable amount of time, our approach delivers desirable outcomes within 5 seconds. This is illustrated in the first example, where FPI successfully transforms a cardboard robot into a cardboard cat. In contrast, Null-Text does not alter the shape of the robot, and EDICT changes it to a cat, but the result does not appear plausible. In the third example, where the task is to change a suitcase to a table, other approaches fail, while ours subtly changes the suitcase into a table by creating a gap between the sides.
Figures 10 and 11 provide further comparisons for real-image editing. In this context, Null-Text and EDICT were executed until convergence, while FPI underwent five iterations per diffusion step. These figures illustrate how alternative approaches may struggle to preserve the original structure or overlook crucial components specified in the target prompt. For example, in the first row of Figure 10, our approach transforms food into flowers, while other methods fail to do so. In the second row, our approach maintains the original placement of the hydrant, whereas Null-Text fails to do so, and EDICT alters the woman’s hair to red. Our approach only changes the hydrant to red without modifying the rest of the image. Figure 11 highlights an example of FPI success (clean water instead of dirty) in the third row, where other alternatives totally fail. The last row of Figure 10 depicts a failure of our method, where it could not transform the cake into a hat as requested.
Figure 12 shows the outcomes of various edits on the same image, providing additional confirmation that FPI selectively modifies only the pertinent parts in the image while maintaining the original structure.
7 Restricting Execution Time for Baselines
In this section, we outline additional implementation details regarding the execution time of Null-Text [13] and EDICT [22].
For Null-Text, we constrained the null optimization iterations during each diffusion to achieve the desired convergence limit. Specifically, for a 5-second duration, we set the iterations to 2, and for 25 seconds, we increased it to 10 optimization steps, following the recommendations of the authors.
Concerning EDICT, the number of diffusion steps during generation was restricted to meet the desired time limit as specified in the paper. For a 5-second duration, we ran EDICT for 10 steps, and for a 25-second duration, we extended it to 35 diffusion steps, out of the full 50 for complete convergence.
8 Modern Approaches for Solving Implicit Functions
In our paper, we utilized fixed-point iterations to solve the implicit function. While more advanced methods like Newton-Raphson and Conjugate Gradient [14] are available for this purpose, we found them impractical for the following reasons: Newton-Raphson exhibited longer convergence times due to gradient computation, and Conjugate Gradients required finding gradients for latents with a dimension of $4\times 64\times 64\approx 16K$, making it impractical.
9 The Effect of Prompt-Aware Adjustment
Figure 13 presents qualitative results demonstrating the impact of prompt-aware adjustment on image inversion and reconstruction, as proposed in Section 3.3 of the main paper. These results provide qualitative justification for the findings presented in the paper, highlighting that, without prompt-aware adjustment, reconstructed images differ from the original.
10 Rare Concept Generation
@@
Figure 14 displays results for the rare-concept generation task introduced in [19, 18]. The objective is to enable the diffusion model to generate concepts rarely seen in its training set by using a few images of that concept. Both methods in [19, 18] utilize diffusion inversion for this purpose. In the main paper, we presented experiments showcasing the impact of our new inversion process on the outcomes of [19, 18].
@@
Here, we provide qualitative results based on NAO [18] centroid computation when utilizing FPI, contrasting them with those obtained through DDIM inversions. The illustrations demonstrate that FPI is capable of identifying high-quality centroids with correct semantic content compared to centroids found by DDIM.

### v2

@@ Appendix B Seed Exploration & Rare concept generation
Fig. S 1: Generating rare concepts based on a few examples with the method of [28] that heavily depends on the diffusion-inversion quality. In our comparison, we evaluate RNRI alongside DDIM (refer to the discussion in the main paper). DDIM frequently struggles to produce a realistic representation of certain objects (such as Oxygen-mask or Patas Monkey) or an accurate depiction of specific concepts (like Tiger-cat or Pay-phone). However, the use of RNRI rectifies these issues in the results. For a detailed quantitative comparison, please refer to the main paper. See the main paper for a quantitative comparison.
@@
Figure S1 further displays results for the rare-concept generation task introduced in [29, 28]. The objective is to enable the diffusion model to generate concepts rarely seen in its training set by using a few images of that concept. Both methods in [29, 28] utilize diffusion inversion for this purpose. In the main paper, we presented experiments showcasing the impact of our new inversion process on the outcomes of [29, 28].
Fig. S 2: The influence of $\lambda$ on reconstruction performance. We observe that employing no regularization ($\lambda=0$) leads to poor reconstruction, while $\lambda=0.1$ typically yields the highest reconstruction accuracy. As $\lambda$ increases, reconstruction accuracy declines, possibly because assigning excessive weight to the prior undermines the root-finding objective.
Fig. S 3: Qualitative comparison for image reconstruction. While DDIM reconstructed images are likely far from the original one, reconstruction from RenNoise still contains inconsistencies (the table at the 1st, lion image distortion or the texture in the teddy bear). RNRI shows nearly exact reconstruction. The accuracy in Reconstruction is later reflected in downstream tasks e.g. editing or long-tail generation. All methods were executed until they achieved convergence.
Fig. S 4: Qualitative comparison for image editing: Illustrates how alternative approaches may struggle to preserve the original structure or overlook crucial components specified in the target prompt, while RNRI succeeds in editing the image properly.
Fig. S 5: Various editing with same input: Note the RNRI capability in both subtle and extensive changes as one would expect from the particular prompt change.
Fig. S 6: Effect of prompt-aware adjustment on image inversion. Here we show the strong impact of our prompt-aware adjustment on the reconstruction results. Without the adjustment, the results are often unrealistic or far from the original image.
@@
We provide qualitative results based on NAO [28] centroid computation when utilizing RNRI, contrasting them with those obtained through DDIM inversions. The illustrations demonstrate that RNRI is capable of identifying high-quality centroids with correct semantic content compared to centroids found by DDIM.
Appendix C The Effect of the prior
In this section, we examine the impact of incorporating our regularization term in the objective function $\mathcal{F}$ in Eq. (7). We measure the effect in two aspects: 1) The remainder for the full objective function $\mathcal{F}$ in Eq. (7) at the predicted solution and the original residual (also known as fidelity term), $r(z_{t})=z_{t}-f(z_{t})$ in Eq. (6). Note that $r(z_{t})$ indicates how accurately the DDIM equation is satisfied. Monitoring the residual $r$ as shown in Figure S2(a) reveals that incorporating our prior not only aids in finding the correct solution (as shown by PSNR results) but also accelerates convergence.
Appendix D Prior trade-off parameter $\lambda$
Figure S2(b) shows the impact of the weight parameter $\lambda$ on the reconstruction accuracy. We observe that using no regularization ($\lambda=0$) results in poor reconstruction, whereas setting $\lambda=0.1$ achieves the highest reconstruction accuracy, underscoring the importance of the prior in the objective function. There is a gradual decrease with larger $\lambda$ values, where at $\lambda=1$ the PSNR starts to decline faster, likely because placing too much emphasis on the prior compromises the residual root-finding objective.
Appendix E Additional Qualitative Results
In this section, we present additional qualitative comparisons involving RNRI and baseline methods. Figure 3 provides insights into inversion-reconstruction, showcasing that RNRI reaches a better reconstruction quality compared to DDIM and ReNoise [10].
Figures 4 provide further comparisons for real-image editing. These figures illustrate how alternative approaches may struggle to preserve the original structure or overlook crucial components specified in the target prompt. Looking at the Latent Diffusion results, in the first row of Figure S4, RNRI correctly transforms food into flowers, while other methods fail to do so. Row six shows an example of RNRI success (clean water instead of dirty), where other alternatives totally fail. The last row of Figure S4 depicts a failure case of our method, where it could not transform the cake into a hat as requested. Editing for Latent Consistency Models (LCM) is more challenging due to the extremely low number of time steps (aimed for speed-up). While ReNoise and RNRI achieve success in the first two rows, rows 3-6 demonstrate instances where ReNoise fails in comparison to RNRI. The last row illustrates a scenario where both methods fail.
Figure S5 shows the outcomes of various edits on the same image, providing additional confirmation that inversion with RNRI yields modification of the pertinent parts in the image while maintaining the original structure.
Appendix F The Effect of Prompt-Aware Adjustment
Figure S6 presents qualitative results demonstrating the impact of prompt-aware adjustment on image inversion and reconstruction, as proposed in Section 4.2 of the main paper. These results provide qualitative justification for the findings presented in the paper, highlighting that, without prompt-aware adjustment, reconstructed images differ from the original.
Fig. S 7: Failure cases: RNRI fails to converge where the prompt and image do not align.
Appendix G RNRI Convergence
Here we provide information about an analysis to find failure cases for convergence We ran inversion and reconstruction, in scale, on COCO2017 (118k caption-image pairs) and found that in 95.4% of cases, RNRI successfully converged to a solution. Specifically, residuals went down from $\sim 1$ (for DDIM) to $<10^{-4}$ indicating convergence, and the PSNR was $>25.7$ indicating good solutions. The remaining $4.6\%$ were samples with incorrect captions, see Fig. S7. These results imply that lack of convergence may indicate text and image miss-alignment which we consider for future work.